OpenStudy (khally92):

given two vectors u and v such that u.v=|u|.|v|. explain what you know about the two vectors? its a vector question under cross product and dot product. thanks

OpenStudy (turingtest):

in general\[\vec u\cdot\vec v=\|\vec u\|\|\vec v\|\cos\theta\]

OpenStudy (turingtest):

now the fact that in this case\[\vec u\cdot\vec v=\|\vec u\|\|\vec v\|\]what does this imply about the value of \(\cos\theta\) in this case?

OpenStudy (khally92):

means titt is 0. because cos tita is 1.

OpenStudy (turingtest):

right, so what does that tell us about the two vectors given that the angle between them is zero?

OpenStudy (khally92):

perpendicular?

OpenStudy (turingtest):

what is the definition of two perpendicular lines?

OpenStudy (khally92):

ummm honestly i do not know?

OpenStudy (turingtest):

that they meet at 90 degrees

OpenStudy (turingtest):

and what did wee say the angle between these two vectors was?

OpenStudy (khally92):

the angle between the two vectos is 0.

OpenStudy (turingtest):

so they aint perpendicular then, right? so try to think what it means for them to meet at 0 degrees

OpenStudy (khally92):

vector with its self magnitude squared.